OpenStudy (anonymous):

I do not understand this at all. Please help i do not want answers i want to understand it. The question is The number of bushels of corn per acre in the United States varies directly as the number of acres of corn planted. If an average of 20 acres yields 1150 bushels of corn, about how many bushels of corn are yielded each year by 72.5 acres. A. 317 bushels B. 1287 bushels C. 4169 bushels D. 8375 bushels

OpenStudy (anonymous):

Let me see what I can do.

OpenStudy (anonymous):

ok Thank You!!

OpenStudy (anonymous):

I would probably say C.4169. What I did was I knew that there was 1150 bushels of corn per 20 acres. So If we divide 1150 by 20, we should get the amount of bushels of corn per acre. So 1150/20 is 57.5 So if 57.5 is bushels of corn per acre, we have to multiply it by 72.5 And 57.5*72.5= 4168.75 Which is closest to C. 4169 And I hope you can forgive me if this is wrong. lol

jimthompson5910 (jim_thompson5910):

you're correct @iPlayin1080p60fps `The number of bushels of corn per acre in the United States varies directly as the number of acres of corn planted` means that `y = k*x` where y = number of bushels of corn x = number of acres planted so if ` 20 acres yields 1150 bushels of corn` then we can use this to find k y = k*x 1150 = k*20 1150/20 = k*20/20 57.5 = k k = 57.5 So the equation is y = 57.5*x Now use x = 72.5 to find y y = 57.5*x y = 57.5*72.5 y = 4,168.75 which rounds to 4169 bushels of corn

OpenStudy (anonymous):

@jim_thompson5910 Thank you for clarifying that! I wasn't sure if I was right. :)

OpenStudy (anonymous):

Thank you sooooooooooooo much!! It makes so much more sense.